Razor:

Okay, I think I already did the- No I didn't, the answer for the first one is 1/2 I think or that's part of the answer, TSO helped me get the finale outcome but now I need to put it in slope intercept form.

Razor:

1 attachment
dude:

Slope intercept is y=mx+b (we know this pls) (There are a couple ways to solve this) We know the slope is 1/2 So \(\large y=\frac12x+b\) In order to find b, we just need to substitute one of the points they gave us

Razor:

one of them? So I can just pick any o.o

dude:

Yes

dude:

Either (8,9) or (-2,4)

Razor:

Let's do uhm... (-2,4) I guess

dude:

Okay so we substitute the x and y coordinate into its corresponding variable x= -2 y= 4 \(\large 4=\frac12(-2)+b\) Now we can solve for b

Razor:

Alright, let me process LOL

dude:

D: The point is in (x,y) form, right? Just substitute the x coordinate in for x and y the for the y (-2,4) ( x, y)

Razor:

4 = 1/2(-2) + 4 ??????????????????????????????????????????????

dude:

No no b is what we are solving for \(\large 4=\frac12(-2)+b\) Do \(\frac12\times \)-2

Razor:

-1

dude:

Yes so we have \(4=-1+b\) We are trying to isolate b What is the opposite of subtracting 1?

Razor:

pSh if I get this wrong I'll be disappointed in myself adding

dude:

Good good c:

dude:

Now add 1 on both sides

Razor:

Oh YEAH cuz we do that, yeah yeah ok

Razor:

5 and 1, if I added on the right sides

dude:



Razor:

WOT

dude:

D: Add 1 on both sides, you are trying to get rid of the -1 -1+1=0 4+1 =5 So, 5 = b

Razor:

It was blank

dude:

Oh rip

dude:

1 attachment
Razor:

Oh I said 1 OOP

dude:

Its okay xD

Razor:

Boom, 5 and 0 is what I MEANT so that means we do like... (5,0).........or

dude:

Well uh 5 and b, i had written it down already but yes Now that we have b, we just put it into the equation we made \(\large y=\frac12x+b\)

Razor:

uhm.... \[4=\frac{ 1}{ 2 }(5) + 1\]

Razor:

I forgot where the five goes

dude:

No no we're just putting in the b value Remember from our last part 5=b

1 attachment
dude:

1 attachment
Razor:

le sigh

dude:

I didn't explain it well, that's on me

Razor:

I'm just me as usual, UHM, so b = 5 so it's gonna be \[4=\frac{ 1 }{ 2 }x + 5\]

dude:

x and y are just iteself now We only substituted the point just to get b, but now that we know what it is we dont need them there \(\large y=\frac12x+5\) is your answer

dude:

Recap: Slope intercept is y=mx+b We know the slope is 1/2 [from Q 1] \(\bf \large y=\frac12x+b\) In order to find b, we just need to substitute one of the points ================================================ We chose (-2,4) Okay so we substitute the x and y coordinate into its corresponding variable x= -2 y= 4 \(\large 4=\frac12(−2)+b\) Now we can solve for b [First multiply \(\frac12\) and-2] \(\large 4=-1+b\) [Now add 1 on both sides to get b by itself] \(5=b\) ================================================ Now that we know what b is, we substitute that into our original equation \(\large y=\frac12x+b\) \(\large \boxed{y=\frac12x+5}\)

Razor:

I was like; ohhhhh frickkkkk I have to read all disssss?!: And then I look up there and it says "Recap" I'm like PHEW! But ye basically that's it, so the answer-- does or doesn't end at y = 1/2x + 5 ?

Razor:

I think it does..

dude:

/makes recap 50x larger Yes that is the answer, we're done with the Q xD

Razor:

xD So the ANSWER for question 1 is 1/2 and the ANSWER for question 2 is y = 1/2x + 5 basically

dude:

Uhh for #1, TSO helped you get the answer We only did #2

Razor:

Yep yep, got it tho I'ma just *closes post and writes the recap so I can TRY to do it myself next time*